# SSD4Rec: A Structured State Space Duality Model for Efficient Sequential Recommendation

### v1

@@ 3.4. Ablation Study
From the ablation results in Table 6, we can see that each component in our approach contributes to the overall performance since eliminating any of them would result in performance degradation. Moreover, the results demonstrate that the integration of learnable positional embeddings does not improve the performance, due to the recurrent nature of the Mamba block. Finally, stacking Mamba layers yields performance improvements on the ML1M dataset with larger, long-sequence user interactions, but it would lead to a linear increase in the model size. Considering the increased complexity resulting from the stacked layers, we have designated the 2-layer SSD4Rec as our default model for performance evaluations.
@@ 3.5. Hyper-parameter Analysis
In SSD4Rec, we introduce two critical hyper-parameters, namely the weighted indicator of backward processing $\beta$ and the max length of item sequences $L$. Their value sensitivities are evaluated in this section to facilitate the future application of our proposed model.
@@ 3.5.1. Effect of Backward Indicator $\beta$
This hyper-parameter determines the degree to which backward information can propagate through the SSD layer. Figure 4 shows the corresponding performance change of SSD4Rec on NDCG@10. We can find that introducing a small ratio of backward knowledge brings performance improvements. In most cases, the recommendation performance of our proposed method improves when $\beta$ 0.3. The experimental results also reveal that the recommendation performance degrades when the masking ratio $\beta$= 0.5, suggesting excessive backward information should be avoided.
@@
Figure 3. The effect of backward weighted indicator $\beta$ under NDCG@10.
@@
(c) KuaiRand - NDCG@10
(d) KuaiRand - MRR@10
@@ 4.1. Sequential Recommendation
1) Traditional methods employ conventional data mining techniques to obtain the users’ preferences from their historical interactions (Davidson et al., 2010; He et al., 2017; Twardowski, 2016; Liu et al., 2023; Zhou et al., 2022). For example, as a typical Markov chains-based method, TransRec He et al. (2017) introduces a novel unified method to model third-order interactions in large-scale sequential prediction tasks by embedding items into a transition space and leveraging translation vectors to capture complex user-item and item-item relationships. Twardowski (2016) propose a matrix factorization-based sequential RecSys to capture user behavior dependencies without explicit user identification, utilizing user activity within single sessions as sequences of events to provide recommendations.
2) Deep learning-based methods leverage the powerful representation capabilities of deep neural networks to model the users’ preferences based on the user-item interaction history. For example, Tang and Wang (2018) propose a Convolutional Sequence Embedding Recommendation Model for top-N sequential recommendation by embedding users’ interaction sequence into an ”image” into the time and latent space and using convolutional filters to capture the sequential patterns. Zhang et al. (2019) utilize the self-attention mechanisms to estimate the weights for each item within the users’ historical behavior to model the users’ preferences for next-item recommendation. Chen et al. (2018) systematically investigates the impact of temporal information in sequential recommendations by uncovering and incorporating two fundamental temporal patterns of user behaviors, ‘absolute time patterns’ and ‘relative time patterns’, into the memory-augmented neural networks to model user dynamic preferences effectively.
@@ 4.2. Structured State Space Models (Mamba)
Due to Mamba’s powerful sequential modeling capabilities and computational efficiency, numerous studies have applied it across various domains in real-life applications (Qu et al., 2024b). For instance, Lieber et al. (2024) proposes a cutting-edge large language model called Jamba that combines Transformer and Mamba layers in a hybrid mixture-of-experts architecture, resulting in a high-performance, memory-efficient model capable of handling long-context sequences effectively. Ruan and Xiang (2024) introduce Vision Mamba UNet (VM-UNet), a novel U-shape architecture model for medical image segmentation that leverages State Space Models (SSMs) to address the limitations of CNNs and Transformers by incorporating a Visual State Space (VSS) block to capture extensive contextual information for medical image segmentation. Xing et al. (2024) a pioneering 3D medical image Segmentation Mamba model that leverages the efficiency of State Space Models (SSM) to effectively capture long-range dependencies within whole-volume features at various scales, outperforming Transformer-based methods in whole-volume feature modeling even with high-resolution volume features and achieving superior processing speed and efficacy in medical image segmentation tasks. For more details on existing works, please refer to a comprehensive survey of mamba (Qu et al., 2024b).
@@ 5. Conclusion
While existing deep learning-based sequential recommendation methods achieve promising prediction performance, they fail to effectively and efficiently capture user preferences with long-tail interactions. To tackle these challenges, we propose a novel framework, named SSD4Rec, based on Mamba architecture for sequential recommendation. To be specific, the proposed SSD4Rec consists of two key modules, i.e., a novel strategy for batch processing of variable-length item sequences and a Bi-SSD block for efficient user behavior sequence modeling. Specifically, using the bidirectional structure enables a comprehensive information understanding, and the SSD layer equips the model with attention-like learning capabilities and superior computational efficiency. Through comprehensive experiments on four distinct datasets, we revealed that our model could achieve state-of-the-art sequential recommendation performance while also exhibiting the capacity to achieve fast training and inference.

### v2

@@ 3.4. Ablation Study
  • w/o mask: Remove the mask operation on the model input, i.e., all inputs remain in their original state.
@@
From the ablation results in Table 7, we can see that each component in our approach contributes to the overall performance since eliminating any of them would result in performance degradation. Moreover, the results demonstrate that the integration of learnable positional embeddings does not improve performance due to the recurrent nature of the Mamba block. Finally, stacking Mamba layers yields performance improvements on the ML1M dataset with larger, long-sequence user interactions, although it would lead to a linear increase in the model size. To this end, as shown in Table 3, we have designated a 2-layer SSD4Rec model for the ML1M and KuaiRand datasets and a 1-layer SSD4Rec model for the Beauty and Games datasets, respectively.
@@ 3.5. Hyper-parameter Analysis
In SSD4Rec, we introduce three critical hyper-parameters, namely the weighted indicator of backward processing $\beta$, the mask ratio indicator $\rho$ and the max length of item sequences $L$. Their value sensitivities are evaluated in this section to facilitate the future application of our proposed model.
@@ 3.5.1. Effect of Backward Indicator $\beta$
This hyper-parameter determines the degree to which backward information can propagate through the SSD layer. Figure 3 shows the corresponding performance change of SSD4Rec on NDCG@10, MRR@10, HR@10, separately. We can find that introducing a small ratio of backward knowledge brings performance improvements. In most cases, the recommendation performance of our proposed method improves when $\beta<0.3$. The experimental results also reveal that the recommendation performance degrades when the backward indicator $\beta\geq 0.5$, suggesting excessive backward information should be avoided. In most cases, the model demonstrates optimal performance when $\beta$ is set to 0.1. Therefore, we designate $\beta=0.1$ as the default value for our proposed method, SSD4Rec.
@@
(e) ML1M - MRR@10
(f) Beauty - MRR@10
(g) Games - MRR@10
(h) KuaiRand - MRR@10
(i) ML1M - HR@10
(j) Beauty - HR@10
(k) Games - HR@10
(l) KuaiRand - HR@10
Figure 3. The effect of backward weighted indicator $\beta$ under NDCG@10, MRR@10, HR@10, separately.
3.5.2. Effect of Mask Ratio indicator $\rho$
This hyperparameter determines the fraction of user-item interaction sequences to be masked. Figure 4 show the corresponding performance change of SSD4Rec on NDCG@10, MRR@10, HR@10 over $\rho$, separately. We can find that introducing a small mask ratio brings performance improvements on most metrics. In most cases, the recommendation performance of our proposed method improves when $\rho\leq 0.2$. The experimental results also reveal that the recommendation performance degrades as the maski ratio increases, especially when $\rho\geq 0.4$, suggesting excessive mask ratio should be avoided. Thus, the ideal mask ratio should take into account the balance between increasing sample diversity to enhance model robustness and preventing excessive information from being masked. Across the four datasets, namely ML1M, Beauty, Games, and Kuairand, the optimal values $\rho_{best}$ vary: specifically, they are 0.1, 0.2, 0.1, and 0.2, respectively.
(a) ML1M - NDCG@10
(b) Beauty - NDCG@10
(c) Games - NDCG@10
(d) KuaiRand - NDCG@10
(e) ML1M - MRR@10
(f) Beauty - MRR@10
(g) Games - MRR@10
(h) KuaiRand - MRR@10
(i) ML1M - HR@10
(j) Beauty - HR@10
(k) Games - HR@10
(l) KuaiRand - HR@10
Figure 4. The effect of mask ratio indicator $\rho$ under NDCG@10, MRR@10, HR@10, separately.
3.5.3. Effect of Max Sequence Length $L$
@@
(c) ML1M - HR@10
(d) KuaiRand - NDCG@10
(e) KuaiRand - MRR@10
(f) KuaiRand - HR@10
@@ 4.1. Sequential Recommendation
1) Traditional methods employ conventional data mining techniques to obtain the users’ preferences from their historical interactions (Davidson et al., 2010; He et al., 2017; Twardowski, 2016; Liu et al., 2023; Zhou et al., 2022). For example, as a typical Markov chains-based method, TransRec He et al. (2017) introduces a novel unified method to model third-order interactions in large-scale sequential prediction tasks by embedding items into a transition space and leveraging translation vectors to capture complex user-item and item-item relationships. Moreover, He and McAuley (2016) integrates similarity-based approaches with Markov Chains to tackle the sparsity challenges in sequential recommendation tasks. Twardowski (2016) propose a matrix factorization-based sequential RecSys to capture user behavior dependencies without explicit user identification, utilizing user activity within single sessions as sequences of events to provide recommendations.
2) Deep learning-based methods leverage the powerful representation capabilities of deep neural networks to model the users’ preferences based on the user-item interaction history. For example, Tang and Wang (2018) propose a Convolutional Sequence Embedding Recommendation Model for top-N sequential recommendation by embedding users’ interaction sequence into an ”image” into the time and latent space and using convolutional filters to capture the sequential patterns. Zhang et al. (2019) utilize the self-attention mechanisms to estimate the weights for each item within the users’ historical behavior to model the users’ preferences for a next-item recommendation. Chen et al. (2018) systematically investigates the impact of temporal information in sequential recommendations by uncovering and incorporating two fundamental temporal patterns of user behaviors, ‘absolute time patterns’ and ‘relative time patterns’, into the memory-augmented neural networks to model user dynamic preferences effectively. Tang et al. (2019) believes that various model architectures excel at capturing different time horizons; some are effective for short-term dependence, while others are better suited for long-term dependence. They propose a hybrid model that incorporates a learned gating mechanism capable of combining different models to effectively manage both short-term and long-term dependencies. Despite effectiveness, most deep learning-based methods still have an intrinsic limitation on capturing long-range dependencies in user-item interactions: RNN-, CNN-, and GNN-based methods exhibit inferior performance stemming from memory compression and limited receptive field, while Attention-based methods encounter time-consuming training and inference due to their quadratic complexity with respect to sequence length.
@@ 4.2. Structured State Space Models (Mamba)
Due to Mamba’s powerful sequential modeling capabilities and computational efficiency, numerous studies have applied it across various domains in real-life applications (Qu et al., 2024b). For instance, Lieber et al. (2024) proposes a cutting-edge large language model called Jamba that combines Transformer and Mamba layers in a hybrid mixture-of-experts architecture, resulting in a high-performance, memory-efficient model capable of handling long-context sequences effectively. Ruan and Xiang (2024) introduces Vision Mamba UNet (VM-UNet), a novel U-shape architecture model for medical image segmentation that leverages State Space Models (SSMs) to address the limitations of CNNs and Transformers by incorporating a Visual State Space (VSS) block to capture extensive contextual information for medical image segmentation. Xing et al. (2024) proposes a pioneering 3D medical image Segmentation Mamba model that leverages the efficiency of State Space Models (SSM) to effectively capture long-range dependencies within whole-volume features at various scales, outperforming Transformer-based methods in whole-volume feature modeling even with high-resolution volume features and achieving superior processing speed and efficacy in medical image segmentation tasks. Hwang et al. (2024) studies a unifying matrix mixer framework for sequence mixers, which includes a wide range of well-known sequence models, such as Transformers and State Space Models (SSMs). Guided by this framework, they propose Hydra, a natural bidirectional extension of Mamba that corresponds to a quasiseparable matrix mixer. Hydra significantly outperforms transformers in non-causal problems. Motivated by this success, (Liu et al., 2024b) develops a Mamba-based model called SIGMA to tackle the challenges of context modeling and short sequence modeling through the introduction of a Partially Flipped Mamba structure, a Dense Selective Gate, and a Feature Extract GRU. These innovations effectively overcome the limitations of the Mamba model, including its unidirectional structure and instability in state estimation for short sequences. For more details on existing works, please refer to a comprehensive survey of mamba (Qu et al., 2024b). Different from existing Mamba-based recommendation approaches, our study presents an innovative Masked Bidirectional Structured Space State Model, exploring the potential of a pure-Mamba model in sequential recommendation.
@@ 5. Conclusion
While existing deep learning-based sequential recommendation methods achieve promising prediction performance, they fail to effectively and efficiently capture user preferences with long-tail interactions. To tackle these challenges, we propose a novel framework, named SSD4Rec, based on Mamba architecture for sequential recommendation. To be specific, the proposed SSD4Rec consists of two key modules, i.e., a novel strategy for batch processing of masked variable-length item sequences and a Bi-SSD Layer for efficient user behavior sequence modeling. Specifically, using the bidirectional structure enables a comprehensive information understanding, and the SSD layer equips the model with attention-like learning capabilities and superior computational efficiency. Through comprehensive experiments on four distinct datasets, we revealed that our model could achieve state-of-the-art sequential recommendation performance while also exhibiting the capacity to achieve fast training and inference.